# Count Every Rotation and Every Rotation Counts: Exploring Drone Dynamics via Propeller Sensing

### v1

@@ 1. Introduction
The drone-driven low-altitude economy, anticipated to reach a \$1 trillion valuation by 2040 (Daily, 2024; Zhang et al., 2025; Chen* et al., 2024), is reshaping industries through transformative services like instant delivery (Chen* et al., 2024; Sharp et al., 2022; Zhao et al., 2024), precision industrial monitoring (Ben Ali et al., 2020; Xia et al., 2023; Ahmad et al., 2023), and emergency response operations (Schedl et al., 2021; Zhang et al., 2023; Sie et al., 2023). Central to this rapidly expanding industry is the critical landing phase (Fig.1), during which ground platforms must rapidly identify drones descending from below 50 meters and guide them to precise touchdown at designated locations (He et al., 2023; Sun et al., 2022).
Current contactless, ground-based localization systems for drone landing primarily install cameras (Zhu et al., 2018; Petrides et al., 2017), acoustic sensors (Bannis et al., 2020; Wang et al., 2022; Bai et al., 2022), and radars (Drozdowicz et al., 2016; Shin et al., 2016; Hoffmann et al., 2016) at the center or along the edges of drone landing pads. These methods excel in low-speed, stable flights but face challenges in complex scenarios involving dynamic accelerations. Such performance degradation stems from the sensors’ limited spatio-temporal resolution (Wang et al., 2025d). For instance, radar often perceives drones merely as point objects and fails to accurately capture 6-degree-of-freedom rotational movements critical for precise landing adjustments (Herschfelt et al., 2017). Similarly, cameras operating at 20-60 FPS frequently may have difficulty profiling high-speed actions that occur during frame intervals (Gehrig and Scaramuzza, 2024; Jiang et al., 2021; Jain et al., 2019).
To overcome these limitations, we expand our perspective beyond individual sensors or algorithms to encompass the full drone propulsion and flight control system, as shown in Fig.2. Given a user-desired or planner-calculated flight trajectory (i.e., a continuous location, linear/angular velocity, and acceleration over time), the drone’s position and attitude controllers produce internal flight commands [^1] Commercial flight controllers such as PX4 (PX4, 2024) standardize drone flight commands in formats like hovering, climbing/descending a specific number of meters, and adjusting pitch/yaw/roll by predetermined degrees (DJI, 2016). at every instant (Mahony et al., 2012). These commands guide each propeller’s motor controller to adjust blade speeds, altering forces and torques, eventually adjusting the drone’s external flight status (i.e., velocity, position, and orientation).
@@
However, challenges still arise as drone propellers can rotate at 2,000 to 10,000 revolutions per minute (RPM) in during typical flights (Deters et al., 2017). Traditional sensors, like cameras, radars, and acoustics, struggle with accurate rotational speed estimation due to limited spatio-temporal resolution (Abdelnasser et al., 2023; Kodukula et al., 2023, 2022; Garg et al., 2021). While sophisticated instruments like motion capture systems and laser velocimeters can measure such speeds, they necessitate deploying markers on propellers (Lin et al., 2022), complicating robust, active, and contactless drone sensing systems.
Nowadays, event cameras, with their ultra-high-frequency and microsecond-level sampling, have revolutionized mobile applications such as high-speed motion tracking (Hao et al., 2023), SLAM (Rebecq et al., 2016), and drone obstacle avoidance (Xu et al., 2022). Unlike conventional cameras, event cameras asynchronously capture pixel-level intensity changes without global exposure times, eliminating motion blur even at extreme velocities (Wang et al., 2023; Xu et al., 2023b). Their unique capabilities place them at the forefront of sensor technology, providing an innovative opportunity for real-time, high-speed propeller rotational speed estimation.
@@
Making EventPro a practical system in the wild presents several challenges that are addressed in this work: $(i)$ how to ensure the efficiency and accuracy of propeller sensing given the immense data output of event cameras (e.g., millions of events per second) and their acute sensitivity to environmental variations, such as noise from drone vibrations and changes in lighting (Gallego et al., 2020); $(ii)$ how to infer both internal control commands and external flight status from the estimated propeller rotational speed. This difficulty arises from the non-linear relationship between control commands and propeller speeds (i.e., reflecting generated force and torques) (Mahony et al., 2012). The design and implementation of EventPro excels in:
$\bullet$ Count Every Rotation (§3). We propose an accurate and efficient event-camera-based propeller rotational speed estimation pipeline, including $(i)$ a distribution-informed hierarchical event preprocessing module to filter out background event noise and segment multiple propellers based on spatiotemporal statistics; $(ii)$ a geometry-instructed event motion compensation algorithm refines standard motion compensation (Gallego et al., 2018) by incorporating propeller geometry information to accurately estimate rotational speed. and $(iii)$ an adaptive event feature representation strategy further enhances efficiency by adjusting and downsampling event batches based on motion consistency and spatiotemporal distribution.
@@
We deployed EventPro with a Prophesee EVK4 HD event camera (Prophesee, 2024) in various experimental testbeds and real-world drone delivery scenarios. Over 20 hours of experiments covered diverse robustness conditions, comparing EventPro’s rotational speed estimation accuracy and latency against two baselines (Zhao et al., 2023; Kolář et al., 2025). Results show that EventPro outperforms both baselines, reducing latency to 3.14msms while achieving a merely 0.23% estimation error, a ¿45% error reduction. EventPro also demonstrates its dual capabilities: it can accurately infer internal drone commands and enhance external flight status estimation. Specifically, it predicts drone flight actions with 96.5% accuracy and then improves drone tracking precision by ¿22% when integrated with existing sensing modalities.
@@ 3. Count Every Rotation
Event cameras are bio-inspired sensors that operate independently and asynchronously for each pixel. Instead of capturing images at a fixed time interval, event cameras record asynchronous changes in pixel brightness, resulting in a stream of events with microsecond resolution (Gallego et al., 2020), as shown in Fig.5a. When a pixel detects a light intensity change exceeding a predefined magnitude, it immediately outputs an event $e_{k}\doteq\left(\mathbf{x}_{k},t_{k},p_{k}\right)$, encoding the triggered position $\mathbf{x}_{k}$, the triggered time $t_{k}$, and polarity $p_{k}$ (+1 for brighter and -1 for darker) of the intensity changes.
@@
$\bullet$ High sensitivity to scene dynamics impairs speed estimation. During operation, the generated noisy events include random background noise (e.g., hot noise, $1/f$ noise(Gallego et al., 2020)) and those triggered by vibrations of the drone’s body and propellers. These noisy events often intermix within the space-time events triggered by propeller rotational motion (Fig.7b), leading to significant interference in estimation.
@@ 3.1. Distribution-informed Hierarchical Event Preprocessing
Figure 5. (a) Event generation principle. (b) The high event-generating speed of two typical commercial drones (Drone-I(DJI, 2024) and Drone-II(Asia, 2023)) even hovering captured by a commercial event camera.
@@ 3.1.1. Temporal Distribution-informed Event Filtering
We transforms the temporal distribution of events into heatmaps to filter noise using statistical metrics. Specifically, for events collected within the time window $t_{o},t_{0}+\delta t]$, they are binned into a 2D histogram based on their coordinates, with each bin representing a specific spatial region (e.g., $5\times 5$ pixels). The characteristics of each bin are usually inconsistent in two aspects: ($i$) Event counts: high-speed rotational motion generates more events than noise. Bins with event counts significantly below the mean (e.g., three times lower) likely contain noise. ($ii$) Event polarities: rotational motion produces bipolar events, whereas background noise tends to be unipolar (Xu et al., 2024). Bins with highly unbalanced distributions of positive and negative events indicate noise.
@@ 3.2. Geometry-instructed Event Motion Compensation
Although preprocessing filters out some noise with distinct features, stubborn residues persist since: ($i$) the sources of noise are multiple (e.g., background activity noise, hot noise, and $1/f$ noise) and their firing time is random (Gallego et al., 2020); and ($ii$) vibrations from the drone’s body and propellers triggers additional events. Consequently, this noisy data is interwoven in the space-time volume, making it difficult to fully filter out and hindering reliable motion estimation.
@@ 3.2.1. Motion Compensation Framework
(c) Optimize the objective function with respect to the parameter $\omega$. Once the objective function is constructed, various optimization tools can be utilized. Note that our algorithm is flexible and not dependent on a specific optimizer. In this work, we apply Brent’s method (Press et al., 1992) to obtain the best model parameters.
@@ 3.2.2. Warping Function
Consider a rotation matrix $\mathrm{R}(t_{k};\omega)$ that describes the 3D rotation motion at $t_{k}$. This matrix can be decomposed into three independent rotations—roll, pitch, and yaw—around each axis, $\mathrm{R}=\mathrm{R}_{X}(\omega_{x})\mathrm{R}_{Y}(\omega_{y})\mathrm{R}_{T}(\omega_{t})$, where $\mathrm{R}_{X},\mathrm{R}_{Y},\mathrm{R}_{T}$ are the rotation matrices of three axes, respectively (mathworld, 2022). For a short period, we assume the rotational angles in the $X$ and $Y$ dimensions remain constant and events are generated mostly by rotational motion around $T$ axis (as shown in Fig.8b). For instantaneous rotational speed estimation over a short time period, we assume the rotational angles about the $X$ and $Y$ axes remain constant and events are generated mostly by rotational motion around $T$ axis (as shown in Fig.8b). Thus the corresponding rotation matrix $\mathrm{R}$ is given by:
@@ 3.3. Adaptive Event Feature Representation
Event cameras generate large data volumes quickly, causing processing delays with increased accumulation time. Variations in device specifications (e.g., resolutions) and target states (e.g., scene dynamics, textures, spatial locations) cause the data volume to vary greatly, from hundreds to tens of thousands of events per millisecond. Current practices often overlook these variations (Xu et al., 2024; Falanga et al., 2020), utilizing fixed batch sizes or fixed time intervals for event accumulation, which leads to information loss or high computational burdens.
@@ 3.3.2. Structure-aware Event Batch Downsampling
With the above dynamic batch size strategy, our estimate algorithm can adapt to various scenarios. However, an event batch may still contain thousands of events, causing computational strain on resource-limited devices. Inspired by existing sampling strategies for 3D point cloud processing (Lin et al., 2024), we design a structure-aware sampling approach called density importance sampling, which effectively retains spatiotemporal information, as depicted in Fig.11b. The basic idea of this method is prioritizing sampling from denser regions of the event point cloud, as these areas are less likely to contain noise and more informative for estimation.
@@ 4.1. Drone Internal Command Inference
We demonstrate the EventPro’ ability to infer the internal flight commands of drones based on the drone’s system-level dynamics. Existing drone monitoring systems primarily focus on detecting or localizing the drone (Abedi and Vasisht, 2022; Chi et al., 2022; Karanam et al., 2018), but lack reliable methods for estimating their intention.
@@
where $d$ is the characteristic length, $m$ is the drone mass, and $J$ is the drone’s mass moment of inertia (Mahony et al., 2012).
Dynamics-informed input. Drone flight commands can be abstracted into six basic motions: hover, yaw, roll, pitch, climb, and descent (Sun et al., 2022). To demonstrate the feasibility of flight command inference, we conducted experiments with a quadcopter (Bitcraze, 2024) that performs various motions for more intuitive analysis. Fig.12 depicts the variation in propeller speed distribution corresponding to changes in drone motion, revealing two key observations: ($i$) during the transition of the drone motion state, there is a noticeable change in propeller speed distribution; and ($ii$) a clear correlation exists between drone motion and its propeller speed distribution. These findings suggest that monitoring the propeller speed over a short period can be a feasible solution for inferring drone flight commands. Consider a drone with $N_{p}$ propellers, we construct the input representation $x_{t}=[\omega_{1,t-T_{h}:t}^{2},\omega_{2,t-T_{h}:t}^{2},\cdots,\omega_{N_{p},t-T_{h}:t}^{2}]$, where $t$ represents the current time step, and $T_{h}$ is the window length that determines the amount of historical data included.
@@ 4.2. Drone External Status Estimation
We demonstrate the feasibility of using EventPro to enhance drone localization. Multirotor autopilots typically rely on onboard sensors or GPS for localization. However, onboard sensors can quickly accumulate errors, and GPS performance degrades in urban canyons at lower altitudes, leading to instability (Wang et al., 2022, 2025e, 2024). Our insight is that the real velocities of a drone’s propellers are fundamental in determining its kinematic state. By integrating the inferred control commands from propeller speed, we can achieve high-precision, low-latency localization that directly enables precise landing assistance. The estimated drone position provides real-time corrective guidance during descent, allowing the drone to adjust its trajectory and accurately align with the target landing pad.
@@
Multi-modal fusion. Next, we utilize the EKF (Thrun, 2002) algorithm to integrate inputs from two sources: GPS, and the predictive positions. EFK is widely regarded as a robust solution in nonlinear systems. It models the drone’s position as a belief $bel(Y^{t})$ with the mean $\mu_{t}$ and the covariance $\Sigma_{t}$. To update these parameters, EFK consisted of two primary steps, including a prediction step and an update step. As shown in Fig.13, in the prediction step, we first utilize the learning-based model to obtain the prediction result $\mu_{t}^{p}$. In the update step, the GPS measurements are used as observations to update the belief. This two-step process is repeated once there are new input data received (Thrun, 2002), and the mean $\mu_{t}^{o}$ of the updated belief is taken as the current estimated position.

### v2

@@ 1. Introduction
The drone-driven low-altitude economy, anticipated to reach a \$1 trillion valuation by 2040 (chinadaily; zhang2025citynavagent; chen2024ddl), is reshaping industries through transformative services like instant delivery (chen2024ddl; sharp2022authentication; zhao2024foes), precision industrial monitoring (ben2020edge; xia2023anemoi; ahmad2023aerotraj), and emergency response operations (schedl2021autonomous; zhang2023rf; sie2023batmobility). Central to this rapidly expanding industry is the critical landing phase (Fig.1), during which ground platforms must rapidly identify drones descending from below 50 meters and guide them to precise touchdown at designated locations (he2023acoustic; sun2022aim).
Current contactless, ground-based localization systems for drone landing primarily install cameras (zhu2018visdrone; petrides2017towards), acoustic sensors (bannis2020bleep; wang2022micnest; bai2022spidr), and radars (drozdowicz201635; shin2016distributed; hoffmann2016micro) at the center or along the edges of drone landing pads. These methods excel in low-speed, stable flights but face challenges in complex scenarios involving dynamic accelerations. Such performance degradation stems from the sensors’ limited spatio-temporal resolution (wang2025mme). For instance, radar often perceives drones merely as point objects and fails to accurately capture 6-degree-of-freedom rotational movements critical for precise landing adjustments (herschfelt2017consumer). Similarly, cameras operating at 20-60 FPS frequently may have difficulty profiling high-speed actions that occur during frame intervals (gehrig2024low; jiang2021flexible; jain2019scaling).
To overcome these limitations, we expand our perspective beyond individual sensors or algorithms to encompass the full drone propulsion and flight control system, as shown in Fig.2. Given a user-desired or planner-calculated flight trajectory (i.e., a continuous location, linear/angular velocity, and acceleration over time), the drone’s position and attitude controllers produce internal flight commands [^1] Commercial flight controllers such as PX4 (px4) standardize drone flight commands in formats like hovering, climbing/descending a specific number of meters, and adjusting pitch/yaw/roll by predetermined degrees (mobilesdk). at every instant (mahony2012multirotor). These commands guide each propeller’s motor controller to adjust blade speeds, altering forces and torques, eventually adjusting the drone’s external flight status (i.e., velocity, position, and orientation).
@@
However, challenges still arise as drone propellers can rotate at 2,000 to 10,000 revolutions per minute (RPM) in during typical flights (deters2017static). Traditional sensors, like cameras, radars, and acoustics, struggle with accurate rotational speed estimation due to limited spatio-temporal resolution (abdelnasser2023multirotor; kodukula2023squint; kodukula2022adaptive; garg2021owlet). While sophisticated instruments like motion capture systems and laser velocimeters can measure such speeds, they necessitate deploying markers on propellers (lin2022wi), complicating robust, active, and contactless drone sensing systems.
Nowadays, event cameras, with their ultra-high-frequency and microsecond-level sampling, have revolutionized mobile applications such as high-speed motion tracking (hao2023eventboost), SLAM (rebecq2016evo), and drone obstacle avoidance (xu2022swarmmap). Unlike conventional cameras, event cameras asynchronously capture pixel-level intensity changes without global exposure times, eliminating motion blur even at extreme velocities (wang2023ev; xu2023visible). Their unique capabilities place them at the forefront of sensor technology, providing an innovative opportunity for real-time, high-speed propeller rotational speed estimation.
@@
Making EventPro a practical system in the wild presents several challenges that are addressed in this work: $(i)$ how to ensure the efficiency and accuracy of propeller sensing given the immense data output of event cameras (e.g., millions of events per second) and their acute sensitivity to environmental variations, such as noise from drone vibrations and changes in lighting (gallego2020event); $(ii)$ how to infer both internal control commands and external flight status from the estimated propeller rotational speed. This difficulty arises from the non-linear relationship between control commands and propeller speeds (i.e., reflecting generated force and torques) (mahony2012multirotor). The design and implementation of EventPro excels in:
$\bullet$ Count Every Rotation (§3). We propose an accurate and efficient event-camera-based propeller rotational speed estimation pipeline, including $(i)$ a distribution-informed hierarchical event preprocessing module to filter out background event noise and segment multiple propellers based on spatiotemporal statistics; $(ii)$ a geometry-instructed event motion compensation algorithm refines standard motion compensation (gallego2018unifying) by incorporating propeller geometry information to accurately estimate rotational speed. and $(iii)$ an adaptive event feature representation strategy further enhances efficiency by adjusting and downsampling event batches based on motion consistency and spatiotemporal distribution.
@@
We deployed EventPro with a Prophesee EVK4 HD event camera (evk4) in various experimental testbeds and real-world drone delivery scenarios. Over 20 hours of experiments covered diverse robustness conditions, comparing EventPro’s rotational speed estimation accuracy and latency against two baselines (zhao2023ev; kolavr2024ee3p3d). Results show that EventPro outperforms both baselines, reducing latency to 3.14msms while achieving a merely 0.23% estimation error, a ¿45% error reduction. EventPro also demonstrates its dual capabilities: it can accurately infer internal drone commands and enhance external flight status estimation. Specifically, it predicts drone flight actions with 96.5% accuracy and then improves drone tracking precision by ¿22% when integrated with existing sensing modalities.
@@ 3. Count Every Rotation
Event cameras are bio-inspired sensors that operate independently and asynchronously for each pixel. Instead of capturing images at a fixed time interval, event cameras record asynchronous changes in pixel brightness, resulting in a stream of events with microsecond resolution (gallego2020event), as shown in Fig.5a. When a pixel detects a light intensity change exceeding a predefined magnitude, it immediately outputs an event $e_{k}\doteq\left(\mathbf{x}_{k},t_{k},p_{k}\right)$, encoding the triggered position $\mathbf{x}_{k}$, the triggered time $t_{k}$, and polarity $p_{k}$ (+1 for brighter and -1 for darker) of the intensity changes.
@@
$\bullet$ High sensitivity to scene dynamics impairs speed estimation. During operation, the generated noisy events include random background noise (e.g., hot noise, $1/f$ noise(gallego2020event)) and those triggered by vibrations of the drone’s body and propellers. These noisy events often intermix within the space-time events triggered by propeller rotational motion (Fig.7b), leading to significant interference in estimation.
@@ 3.1. Distribution-informed Hierarchical Event Preprocessing
Figure 5. (a) Event generation principle. (b) The high event-generating speed of two typical commercial drones (Drone-I(m30t) and Drone-II(meituan)) even hovering captured by a commercial event camera.
@@ 3.1.1. Temporal Distribution-informed Event Filtering
We transforms the temporal distribution of events into heatmaps to filter noise using statistical metrics. Specifically, for events collected within the time window $t_{o},t_{0}+\delta t]$, they are binned into a 2D histogram based on their coordinates, with each bin representing a specific spatial region (e.g., $5\times 5$ pixels). The characteristics of each bin are usually inconsistent in two aspects: ($i$) Event counts: high-speed rotational motion generates more events than noise. Bins with event counts significantly below the mean (e.g., three times lower) likely contain noise. ($ii$) Event polarities: rotational motion produces bipolar events, whereas background noise tends to be unipolar (xu2024seeing). Bins with highly unbalanced distributions of positive and negative events indicate noise.
@@ 3.2. Geometry-instructed Event Motion Compensation
Although preprocessing filters out some noise with distinct features, stubborn residues persist since: ($i$) the sources of noise are multiple (e.g., background activity noise, hot noise, and $1/f$ noise) and their firing time is random (gallego2020event); and ($ii$) vibrations from the drone’s body and propellers triggers additional events. Consequently, this noisy data is interwoven in the space-time volume, making it difficult to fully filter out and hindering reliable motion estimation.
@@ 3.2.1. Motion Compensation Framework
(c) Optimize the objective function with respect to the parameter $\omega$. Once the objective function is constructed, various optimization tools can be utilized. Note that our algorithm is flexible and not dependent on a specific optimizer. In this work, we apply Brent’s method (press1992numerical) to obtain the best model parameters.
@@ 3.2.2. Warping Function
Consider a rotation matrix $\mathrm{R}(t_{k};\omega)$ that describes the 3D rotation motion at $t_{k}$. This matrix can be decomposed into three independent rotations—roll, pitch, and yaw—around each axis, $\mathrm{R}=\mathrm{R}_{X}(\omega_{x})\mathrm{R}_{Y}(\omega_{y})\mathrm{R}_{T}(\omega_{t})$, where $\mathrm{R}_{X},\mathrm{R}_{Y},\mathrm{R}_{T}$ are the rotation matrices of three axes, respectively (euler). For a short period, we assume the rotational angles in the $X$ and $Y$ dimensions remain constant and events are generated mostly by rotational motion around $T$ axis (as shown in Fig.8b). For instantaneous rotational speed estimation over a short time period, we assume the rotational angles about the $X$ and $Y$ axes remain constant and events are generated mostly by rotational motion around $T$ axis (as shown in Fig.8b). Thus the corresponding rotation matrix $\mathrm{R}$ is given by:
@@ 3.3. Adaptive Event Feature Representation
Event cameras generate large data volumes quickly, causing processing delays with increased accumulation time. Variations in device specifications (e.g., resolutions) and target states (e.g., scene dynamics, textures, spatial locations) cause the data volume to vary greatly, from hundreds to tens of thousands of events per millisecond. Current practices often overlook these variations (xu2024seeing; falanga2020dynamic), utilizing fixed batch sizes or fixed time intervals for event accumulation, which leads to information loss or high computational burdens.
@@ 3.3.2. Structure-aware Event Batch Downsampling
With the above dynamic batch size strategy, our estimate algorithm can adapt to various scenarios. However, an event batch may still contain thousands of events, causing computational strain on resource-limited devices. Inspired by existing sampling strategies for 3D point cloud processing (lin2024e2pnet), we design a structure-aware sampling approach called density importance sampling, which effectively retains spatiotemporal information, as depicted in Fig.11b. The basic idea of this method is prioritizing sampling from denser regions of the event point cloud, as these areas are less likely to contain noise and more informative for estimation.
@@ 4.1. Drone Internal Command Inference
We demonstrate the EventPro’ ability to infer the internal flight commands of drones based on the drone’s system-level dynamics. Existing drone monitoring systems primarily focus on detecting or localizing the drone (abedi2022non; chi2022wi; karanam2018magnitude), but lack reliable methods for estimating their intention.
@@
where $d$ is the characteristic length, $m$ is the drone mass, and $J$ is the drone’s mass moment of inertia (mahony2012multirotor).
Dynamics-informed input. Drone flight commands can be abstracted into six basic motions: hover, yaw, roll, pitch, climb, and descent (sun2022aim). To demonstrate the feasibility of flight command inference, we conducted experiments with a quadcopter (crazyfile) that performs various motions for more intuitive analysis. Fig.12 depicts the variation in propeller speed distribution corresponding to changes in drone motion, revealing two key observations: ($i$) during the transition of the drone motion state, there is a noticeable change in propeller speed distribution; and ($ii$) a clear correlation exists between drone motion and its propeller speed distribution. These findings suggest that monitoring the propeller speed over a short period can be a feasible solution for inferring drone flight commands. Consider a drone with $N_{p}$ propellers, we construct the input representation $x_{t}=[\omega_{1,t-T_{h}:t}^{2},\omega_{2,t-T_{h}:t}^{2},\cdots,\omega_{N_{p},t-T_{h}:t}^{2}]$, where $t$ represents the current time step, and $T_{h}$ is the window length that determines the amount of historical data included.
@@ 4.2. Drone External Status Estimation
We demonstrate the feasibility of using EventPro to enhance drone localization. Multirotor autopilots typically rely on onboard sensors or GPS for localization. However, onboard sensors can quickly accumulate errors, and GPS performance degrades in urban canyons at lower altitudes, leading to instability (wang2022micnest; wang2025aerial; wang2024transformloc). Our insight is that the real velocities of a drone’s propellers are fundamental in determining its kinematic state. By integrating the inferred control commands from propeller speed, we can achieve high-precision, low-latency localization that directly enables precise landing assistance. The estimated drone position provides real-time corrective guidance during descent, allowing the drone to adjust its trajectory and accurately align with the target landing pad.
@@
Multi-modal fusion. Next, we utilize the EKF (thrun2002probabilistic) algorithm to integrate inputs from two sources: GPS, and the predictive positions. EFK is widely regarded as a robust solution in nonlinear systems. It models the drone’s position as a belief $bel(Y^{t})$ with the mean $\mu_{t}$ and the covariance $\Sigma_{t}$. To update these parameters, EFK consisted of two primary steps, including a prediction step and an update step. As shown in Fig.13, in the prediction step, we first utilize the learning-based model to obtain the prediction result $\mu_{t}^{p}$. In the update step, the GPS measurements are used as observations to update the belief. This two-step process is repeated once there are new input data received (thrun2002probabilistic), and the mean $\mu_{t}^{o}$ of the updated belief is taken as the current estimated position.